OpenStudy (anonymous):

Trying to remember how to calculate sin^3(pi/4)? Feeling very foolish that it's not coming to mind. I know that the pi/4 value is (sqrt 2)/2. Trying to remember how the cubed sine impacts it. thanks for the help. :)

OpenStudy (amistre64):

do you recall the sin of pi/4?

OpenStudy (anonymous):

\[\sqrt{2}/2\]

OpenStudy (amistre64):

\[\left(\frac {2^{1/2}}{2}\right)^3\] \[\Large\frac {2^{(3*\frac12)}}{2^3}\]

OpenStudy (amistre64):

it just notational, but these are equivalent \[sin^3(x)\to\ [sin(x)]^3\]

OpenStudy (anonymous):

That was what I was thinking but my answer kept coming up different than what I was seeing in the book. I was trying to work it out to see where my error was coming. Looks like it was simple math issues rather than concept problem. Thanks for the help.

OpenStudy (amistre64):

youre welcome